OpenStudy (anonymous):

WILL MEDAL AND FAN!, please help me so i can pass this class Find the x-intercepts of the parabola with vertex (-3,16) and y-intercept (0,-20). Write your answer in this form: (x1,y1),(x2,y2). If necessary, round to the nearest hundredth.

OpenStudy (anonymous):

@AliceCullen help please?

OpenStudy (anonymous):

x = -5 and x = -1

OpenStudy (anonymous):

x-intercepts: -5 and -1

OpenStudy (anonymous):

youre sure???

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Given equation: y=A(x+3)^2+16 using given point (0,-20) to solve for A -20=A(0+3)^2+16 9A=-36 A=-4 Equation of parabola: y=-4(x-h)^2+k Find x-intercepts set y=0 0=-4(x+3)^2+16 4(x+3)^2=16 (x+3)^2=4 x+3±√4=±2 x=-5 or x=-1 Ans: x-intercepts: -5 and -1

OpenStudy (anonymous):

it said incorrect, i think i just typed it in wrong though, could you help me with another please?

OpenStudy (anonymous):

Find the x-intercepts of the parabola with vertex (-1,-108) and y-intercept (0,-105). Write your answer in this form: (x1,y1),(x2,y2). If necessary, round to the nearest hundredth.